Hello,
We are learning about similar triangles.
Here is the question:
If MNPQ is a parallelogram, find RN, RP, and SP.
The answer the book says is that the triangles QSM~PSR~NMR are similar b/c of the Angle Angle Similarity rule (RN=33.3, RP=13.3, SP=20). Which 2 angles are similar?
Thanks.
Photobucket

Answer by rfer(12662)

400+900=1300
ms=36
------------
30/36=50/x
30x=1800
x=60
mr=60
--------------
36/60=20/x
36x=1200
x=33.3
rn=33.3
-------------
sr=60-36
sr=24
--------------
50/60=x/24
60x=1200
x=20
sp=20
-----------
20/30=x/20
30x=400
x=13.3
rp=13.3
-----------
angles q & n are simular